A wheel of diameter 0.3 m is rotating with an angular velocity of 3 rad/s and slowing under a constant acceleration of -3 rad/s2. Calculate everything possible about the motion, starting with, say, its angular velocity after 0.4 s.
Answer #1


w = wo+alfa*t


after t = 0.4
w = 3 - (3*0.4*0.4) = 2.52 rad/s <-answer

angular displacement = theta = wo*t + 0.5*alfa*t^2

theta = 3*0.4 -0.5*3*0.4*0.4 = 0.96 radians <-answer

wf = 0


0 = 3 - 0.3*T

T = 10 s


after T = 10 s thet wheel stops


total displacement = 3*10-0.5*0.3*10*10 = 15 radians = 2.4 revolutions